# Understanding Ethereum Mempool Security under Asymmetric DoS by Symbolic Fuzzing

### v2

@@ 8 Discussions
Defenses: Rule-based mitigation can fix some of our discovered bugs: $XT_{4}$ is fixed by detecting latent overdraft transactions upon replacement. Locking-based $XT_{8}$ is fixed by re-enabling eviction and replacement; to avoid reintroducing eviction vulnerabilities, the fix needs to comply with the security rules against eviction attacks.
Ethical concerns: When evaluating attacks, we only mounted attacks on the testnet and did so with minimal impacts on the tested network. For instance, our attack lasts a short period of time, say no more than $4$ blocks produced. We did not test our attack on the Ethereum mainnet. When reporting bugs, we disclose to the developers the mitigation design tradeoff and the risk of fixing one attacks by enabling other attacks (described above). To prevent introducing new bugs, we did not suggest fixes against locking ($XT_{8}$).
Private and public mempools: In Ethereum, transactions can be sent privately or publicly. A private transaction received on a block builder is directly packed to a block without broadcast to other builders. By contrast, a public transaction received on a node is broadcast to other nodes before it is included in the next block.
On the one hand, whether a mempool stores private or public transactions is orthogonal to the admission/eviction policy the mempool runs. Empirically, we found three open-source Ethereum block builders, named EigenPhi builder, Flashbot builder $v1.11.5-0.2.1$ and bloXroute builder. Specifically, Flashbot and bloXrout builder run the same admission policy as Geth $v1.11.4$, while EigenPhi builder runs the same admission policy as Geth $<v1.11.4$.Therefore, these private mempools are subject to the same vulnerabilities as the public mempools.
On the other hand, a private mempool receives transactions from a portion of the web3 users, while a public mempool receives transactions from all web3 users. Thus, a disabled private mempool creates limited damage by victimizing the portion of users connected to the pool. The damage of a disable public mempool can be amplified by broadcasting the payload transactions across the network and reaching the public mempools of all nodes.
9 Conclusion
This paper presents mpfuzz, the first mempool fuzzer to find asymmetric DoS bugs by symbolically exploring mempool states and optimistically estimating the promisingness an intermediate state is in reaching bug oracles. Running mpfuzz on six major Ethereum clients leads to the discovery of new mempool vulnerabilities, which exhibit a wide variety of sophisticated patterns including stealthy mempool eviction and mempool locking.

### v3

@@ D.1 Experiment on a Single Victim Node
Flashbot builder $v1.11.5$ $XT_{6}$ $94.74\%$ $0.0022$ $11.39$
EigenPhi $XT_{1}$ $99.80\%$ 00 $11.39$
builder $XT_{2}$ $99.42\%$ $0.725$ $11.39$
$XT_{3}$ $93.02\%$ $0.0021$ $11.39$
$XT_{4}$ $99.42\%$ $0.806$ $11.39$
$XT_{6}$ $94.74\%$ $0.0022$ $11.39$
bloXroute builder-ws $XT_{6}$ $94.74\%$ $0.0022$ $11.39$
go-opera $v1.1.3$ $XT_{2}$ $99.12\%$ $0.201$ $11.39$
$XT_{3}$ $92.60\%$ $0.0021$ $11.39$
$XT_{4}$ $99.13\%$ $0.221$ $11.39$
$XT_{6}$ $93.76\%$ $0.0022$ $11.39$
BSC $v1.3.8$ $XT_{6}$ $94.74\%$ $0.0022$ $11.39$
core-geth $v1.12.18$ $XT_{6}$ $94.74\%$ $0.0022$ $11.39$
[…]
OpenEthereum $v3.3.5$ $XT_{4}$ $99.56\%$ $0.233$ $11.39$
$XT_{9}$ $98.36\%$ $0.177$ $18.35$
The evaluation results of all attacks on all clients are in Table 6. On the six Ethereum clients in the public transaction path, the attack success rates are all higher than $84.63\%$, and the attack costs are all lower than $1.172$ Ether per block, which is significantly lower than the baseline.
We also evaluate the attacks found on PBS and Ethereum-like clients. Under the same experimental settings as § D.1, the results show similar success rates and costs with the attacks on Geth (recall these clients are Geth forks). More specifically, Table 6 shows that the success rates are higher than $92.60\%$, and attack costs are lower than $0.806$ Ether per block.